One number is equal to the product of -2 and the difference between another number (n) and 5. The sum of the numbers id 16. What is the second number?
Answer by CubeyThePenguin(3113)

The two numbers are m and n.

m = -2(n-5)
m + n = 16

Plug in the first equation into the second.

-2(n-5)+n=16
-2n+10+n=16
n = -6

The numbers are 22 and -6.